OpenStudy (anonymous):

I guess I'm having issues with fractions... x/2-5=x/7+2

OpenStudy (unklerhaukus):

\[\frac x2-5=\frac x7+2\] right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok bro, the answer is: